OpenStudy (anonymous):

please help with LAB 12

OpenStudy (anonymous):

Given for me for lab is The goal is to create a filter with a resonant frequency ω O =10 kHz±100 Hz and a Q>5, Please help

OpenStudy (anonymous):

i have the values R1=10 ohms, C1=0.1*10^-6, R2=2490 ohms, C2=0.1*10^-6

OpenStudy (anonymous):

Thanks I was typing 1 wrong

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

Not correct for me help